Solve each system by the addition method. Determine whether the equations are independent, dependent, or inconsistent
4x +3y = 2
-12x – 9y = -6

Answer by richwmiller(17219)

multiply first equation by 3
4x+9y=6
-12x-9y=-6
add
-8x=0
x=0
y=2/3
(0,2/3)
for independent etc see
http://mathforum.org/library/drmath/view/62538.html